Calculate the energy released when 50.0g of steam condenses at 100℃.

Use the cooling curve to show what happens when 50.0g steam condenses at 100℃ and then cools to 20℃. Then calculate this energy change.


6. Use the below cooling curve for water for the following questions. 

A. Label the melting/freezing point. 

B. Label the boiling/evaporation point. 

C. At what state is water when the temperature is 80 °c? 

D. At what state is water when the temperature is -20 °C? 

E. What happens to the temperature of water when energy is removed from water at 70°C? 

F. What happens to the temperature of water when energy is removed from water at 100 °C? 

G. Label the points on the plot where each of the following constants are used to determine the calories associated with a change in heat. Hlas, Hvap, Specific Heat (SH or c) 

H. Calculate the energy released when 50.0 g of steam condenses at 100 °C. 

I. Use the cooling curve to show what happens when 50.0 g steam condenses at 100 °C and then cools to 20 °C. Then calculate this energy change. 
Answer #1

For the labeling parts (A, B, G):

140 120 100 80 Temp. 60 40 20 C) 20 Hvap Boiling Point Freezing Point Hfus Heat Removed

For letter C, water is on a liquid state at 80oC as it is between the freezing and boiling point.

For letter D, water is on a solid state at -20oC as it is after the freezing point.

For letter E, temperature drops from 70oC when energy is removed, as there is no change in state.

For letter F, temperature remains the same at 100oC when energy is removed until gas changes to a liquid state.

For letter H:

Heat of vaporization of water is 44 kJ / mol

We turn the 50 grams to moles with the molar mass:

50 g * (1 mol / 18 g) = 2.78 moles

So the energy released will be:

Q = 2.78 * 44 = -122.32 kJ (negative sign as it is heat released)

For letter I, we add the heat change of water from 100 to 20 degrees using the heat capacity of water:

Q = -122.32 kJ + (50g * 4.14 J/goC * (20-100)oC)*(1 kJ / 1000 J)

Q = -122.32 + (-16.56) = -138.88 kJ